OpenStudy (rock_mit182):

Numerical palindromes A palindrome is an integer, such as 45654, that reads the same backward and forward. (a) How many five-digit palindromes are there? (b) How many n-digit palindromes are there?

OpenStudy (rock_mit182):

Please, Anyone ?

OpenStudy (raden):

for 5-digit polindromes, we can form it like ABCBA for the first digit A, we can choice 9 ways for the second digit = 10 ways for the third digit = 10 ways and the rest is only 1 way so, totaly is 9 x 10 x 10 = 900

OpenStudy (raden):

corect me if im wrong, @rational :D

OpenStudy (rock_mit182):

in order ro solve (b) we have ... (n-2) * (n-1) * n * 1 * n * (n-1) *(n-2).... Im not even sure if that makes sense..

OpenStudy (rock_mit182):

IN fact i would say: \[n*(n-1)*(n-2)................1...................(n-1)*(n-2)*n\]

OpenStudy (rock_mit182):

or maybe \[\sum_{i=1}^{m}(n-m+k)\] where k =1 and n at some point has to be equal to m

OpenStudy (rational):

yes looks good to me :) \[ABCBA\] total number of 5 digit palindromes = \(9*10*10\) because the last two digits are determined by first two digits.

OpenStudy (rational):

Similarly if \(n=2k+1\) is odd, then the total number of \(n\) digit palindromes is \[9*10^{k}\]

OpenStudy (rational):

and if \(n=2k\) is even, then the total number of \(n\) digit palindromes is \[9*10^{k-1}\]

OpenStudy (rock_mit182):

k has to be an integer as well i gues..

OpenStudy (rock_mit182):

i dont get it, if k =1 for an even number, what does it mean ? \[9*10^{1-1}= 9*1\]

OpenStudy (rock_mit182):

or if k equals zero for an even number: \[ 9 * 10^{-1}\]

OpenStudy (rational):

n = 2k is an even number no matter what the value of k is

OpenStudy (rational):

n = 2k+1 is an odd number no matter what the value of k is

OpenStudy (rational):

they are the definitions of even and odd numbers

OpenStudy (rock_mit182):

look for example when k is less or equal to 0 we have number like this \[9*10^{-2}\] = \[9*{\frac{1}{10^{2}}} \]

OpenStudy (rock_mit182):

i guess speaking of palindrums this number does not represent at all this concept

OpenStudy (rational):

Notice here "n" represents number of digits, so what do you mean by a number having negative number of digits ?

OpenStudy (rock_mit182):

and that's why i said k has to be equal or greater than 2

OpenStudy (rock_mit182):

;d anyway thanks for helping me out